OpenStudy (anonymous):

how u solve this one @.@ Provide an appropriate response. Find a linear function f(x) with slope -3 such that f(-8)=6

OpenStudy (anonymous):

here, the equation is y-y1=m(x-x1), where m is slope (-3), and x1, y1 are the points that are given to you, namely (-8,6). PLug that into the eqn to get y=6+(-3)(x-(-8)) y= 6-3x-24 y=-3x-18

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

so wat is ur answer then

OpenStudy (anonymous):

i already gave the answer. its y=-3x-18

OpenStudy (anonymous):

kk thax